OpenStudy (anonymous):

find the roots of 2x^2-5x+1=0

OpenStudy (anonymous):

2.28 n 0.22 :P

OpenStudy (anonymous):

the roots of the equation ax^2 + bx +c =0 is \[(-b \pm \sqrt{b ^{2}-4ac})/2a\] solve then

OpenStudy (anonymous):

i plug in the values for a b c in the lower equation correct

OpenStudy (anonymous):

\[-(-5)\pm \sqrt{-5^2-4(2)(1)\div2(2)}\] like that

OpenStudy (jit4won):

−52−4(2)(1) only

OpenStudy (jit4won):

then divide your answer with 2(2)

OpenStudy (anonymous):

\[5\pm \sqrt{25-8}\div4\]

OpenStudy (jit4won):

yes

OpenStudy (anonymous):

\[5\pm \sqrt{18\div4}\]

OpenStudy (jit4won):

Nope 4 is not with 18

OpenStudy (anonymous):

now im confused on what to do next

OpenStudy (jit4won):

Forget 4 and write ur answer

OpenStudy (anonymous):

 not 18